How do you put 4x^2+y^2+24x-4y+36=0 into graphing form?
Answer by ewatrrr(24785)


Hi
4x^2+24x +y^2-4y+36=0
4(x^2+6x) +y^2-4y+36=0
4(x+3)^2 -36+ (y-2)^2 - 4 + 36 = 0 COMPLETING THE SQUARES
4(x+3)^2 + (y-2)^2 = 4
(x+3)^2 + (y-2)^2 /4 = 1
(x+3)^2 + (y-2)^2 /2^2 = 1 Ellipse: C(0,2) Y-AXIS THE MAJOR AXIS